When doing an equation such as...
Negatuve one eigth of a number is -3/4
What am I to do
: When doing an equation such as...
Negatuve one eigth of a number is -3/4
What am I to do

Answer by checkley77(3394)

-1/8*X=-3/4
-X/8=-3/4 NOW CROSS MULTIPLY
-4X=-24
X=-24/-4
X=6 ANSWER.
PROOF:
-6/8=-3/4
-3/4=-3/4